Comparative financial statements for Weller Corporation, a merchandising company, for the year ending December 31 appear below. The company did not issue any new common stock during the year. A total of 700,000 shares of common stock were outstanding. The interest rate on the bond payable was 10%, the income tax rate was 40%, and the dividend per share of common stock was $0.75 last year and $0.40 this year. The market value of the company’s common stock at the end of this year was $27. All of the company’s sales are on account.

Weller Corporation
Comparative Balance Sheet
(dollars in thousands)
This Year Last Year
Assets
Current assets:
Cash $ 1,270 $ 1,340
Accounts receivable, net 9,600 8,200
Inventory 13,400 12,100
Prepaid expenses 750 620
Total current assets 25,020 22,260
Property and equipment:
Land 9,500 9,500
Buildings and equipment, net 53,031 39,485
Total property and equipment 62,531 48,985
Total assets $ 87,551 $ 71,245
Liabilities and Stockholders' Equity
Current liabilities:
Accounts payable $ 20,000 $ 18,300
Accrued liabilities 970 850
Notes payable, short term 240 240
Total current liabilities 21,210 19,390
Long-term liabilities:
Bonds payable 9,000 9,000
Total liabilities 30,210 28,390
Stockholders' equity:
Common stock 700 700
Additional paid-in capital 4,000 4,000
Total paid-in capital 4,700 4,700
Retained earnings 52,641 38,155
Total stockholders' equity 57,341 42,855
Total liabilities and stockholders' equity $ 87,551 $ 71,245

Total liabilities and stockholders equity $87,551 $71,245 Weller Corporation Comparative Income Statement and Reconciliatiorn (dollars in thousands) This Year Last Year Sales Cost of goods sold Gross margin Selling and administrative expenses: $79,210 $66,000 35 70035,000 43,510 31,000 10,600 10,500 7,4006,700 17,200 13,800 900 12,900 5,160 7,740 525 7,215 38,155 30,940 Selling expenses Administrative expenses Total selling and administrative expenses Net operating income Interest expense Net income before taxes Income taxes Net income Dividends to common stockholders Net income added to retained earnings Beginning retained earnings 18,000 25, 510 900 24,610 9,844 14,766 280 14,486 $52,641$38,155 Ending retained earnings Required: Compute the following financial data for this year: ts recsiy able turnover. (Assume that all sales are on account.) (Round your answe 1 Arcounts receiy

Required:

Compute the following financial data for this year:

1. Accounts receivable turnover. (Assume that all sales are on account.) (Round your answer to 2 decimal places.)

2. Average collection period. (Use 365 days in a year. Round your intermediate calculations and final answer to 2 decimal places.)

3. Inventory turnover. (Round your answer to 2 decimal places.)

5. Operating cycle. (Round your intermediate calculations and final answer to 2 decimal places.)

6. Total asset turnover. (Round your answer to 2 decimal places.)

pou. (ose 36s udys Ih a year. Round your intermediat eBook Hint Print References 3. Inventory turnover. (Round your answer to 2 decimal places.) 4. Average sale period. (Use 365 days in a year. Round your intermediate cale 5. Operating cycle. (Round your intermediate calculations and final answer to 6. Total asset turnover. (Round your answer to 2 decimal places.) 1. Accounts receivable turnover 2. Average collection period 3. Inventory turnover 4. Average sale period 5. Operating cycle 6. Total asset turnover days days days Mc Graw < Prev4 of 7
Answer #1

1) Account receivable turnover = Sales/Average receivable = 79210/8900 = 8.90 Times

2) Average collection period = 365/8.9 = 41.01 days

3) Inventory turnover = Cost of goods sold/Average inventory = 35700/12750 = 2.80 Times

4) Average sale period = 365/2.8 = 130.36 days

5) Operating cycle = 130.36+41.01 = 171.37 days

6) Assets turnover = Sales/Average assets = 79210/79398 = 0.99 Times